OpenStudy (anonymous):

8x+3y=-9 what is the slope of this line

OpenStudy (icalibear):

Formula for slope- x1-x2/y1 - y2

OpenStudy (anonymous):

but how do i solve it ?

OpenStudy (icalibear):

Plug in the numbers and solve

OpenStudy (anonymous):

m is slope. so if your equation for slope is y = m x + b

OpenStudy (anonymous):

so the answer is 8

OpenStudy (icalibear):

oh wrong formula my bad its mx+b lol

OpenStudy (anonymous):

You get Y by itself on the left side. 3y=-8x-9 Then you divide everything by 3 because you want y by itself y=-8/3x-3 The equation of the like is y=mx+b where "m" is the slope so the slope is "-8/3"

OpenStudy (anonymous):

k